Questions 1 – 4 refer to prospects X and Y below, as well as the following information
?? = ($0, 0.50; $40, 0.25; $100, 0.25) ?? = ($0, 0.25; $40, 0.50; $60, 0.25)
Mark has utility of wealth given by ?(?) = ?^0.6
1. What is the expected value of prospect X (?V(?))?
2. What is the standard deviation of prospect X (SD(?))? (Round your answer to the nearest cent,
and don’t worry, I’ll include a healthy margin of error so you won’t get this wrong due to
rounding).
3. What is the value of the expected utility of Y for Mark (EU(Y))?
4. How much is Mark willing to spend to acquire Y (so, what is the value of CE(?))? (Hint, if
?(?) = ?^0.6, ?ℎ?n u^−1(?) = ?^1.67).
Answer #1

1.

Expected value of prospect X = 0.5* $0 + 0.25 * $40 + 0.25 * $100

Expected value of prospect X = $35

2.

X P(x) X* P(x) X^2*P(x)
0 0.5 0 0
40 0.25 10 400
100 0.25 25 2500
E(x) 35 2900
Var 1675
Std dev 40.93

Note that,

Variance = X^2 * P(x) - (E(x))^2 = 2900 - 35^2

Variance = 1675

SD = sqrt ( Variance ) = sqrt( 1675) = 40.93

3.

Similar to the above method,

Y P(y) U(y) = y^0.6 U(y)* P(y)
0 0.25 0.00 0
40 0.5 9.15 4.57
60 0.25 11.67 2.92
E(U(y)) 7.49

Hence,

Expected Value = 7.49

4.

Y P(y) U(y) = y^0.6 U(y)^1.67 C(y) =(U(y)^1.67)* P(y)
0 0.25 0.00 0 0.00
40 0.5 9.15 40.33 20.16
60 0.25 11.67 60.53 15.13
E(C(y)) 35.29

Hence,

Value of CE(y)= 35.29